$38,520 is deposited into an account that earns 5.8% interest,which is compounded daily.How much money will be in the account after 8 years?
Answer by addingup(3677)

Daily interest rate: 0.058/365 = 0.0001589
38,520(1.0001589)^365*8 = 38,520*1.59 = 61,246.80
after 8 years there will be 61,246.80 in the account